OpenStudy (anonymous):

Use Lagrange's method to determine the critical points of z = f(x,y) = 2x + 2y subject to x^2 + y^2 + z^2 = 9

OpenStudy (dan815):

soo partial derivative way no good?

OpenStudy (anonymous):

It is:) But this question was in my test this morning and specifically asked to use L's method... Help pls:| hah

OpenStudy (dan815):

ah well i dont know what lagrange is xD, i did see a video titles lagrange method on the MIT lecture site, u want link?

OpenStudy (anonymous):

That would help a lot, thanks!

OpenStudy (dan815):

u cud also try to find the gradient equation and see where its normal is pointed in K direction only

OpenStudy (dan815):

ok gimme sec

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (dan815):

u r welcome :)

OpenStudy (accessdenied):

While I entirely trust the effectiveness of MIT's video lectures (I've used them from time to time!), I just want to include one other thing I find myself often referring to when I see Lagrange Multipliers on OpenStudy (since despite going over the method every time, it still eludes me when I find it again, such as now!) >> http://tutorial.math.lamar.edu/Classes/CalcIII/LagrangeMultipliers.aspx

OpenStudy (anonymous):

I'm interested in this method, could you please put the final solution here so I can check if I did it right? Thanx!

OpenStudy (amistre64):

F =2x + 2y - z, G = L(x^2 + y^2 + z^2 - 9) Fx =2, Gx = L 2x Fy =2, Gy = L 2y Fz = -1, Gz = L 2z equating the F and G parts x = 1/L y = 1/L z = -2/L solve for L in the restraint x^2 + y^2 + z^2 = 9 1/L^2 + 1/L^2 + 4/L^2 = 9 5/L^2 = 9 5/9 = L^2 L = +- sqrt(5)/3